please verify the identity.
2(cos’ x + 1) = 5cos x
Answer #1


solution! 2 ( cos x + 1) = 5 cos x a cos²x + 2 = 5 cosx 2 cos²x - 5 cosx + 2 = 0 2005hp - Acosx- Cosx + 220 & Cux (Ce4 x - 3)